OpenStudy (anonymous):

Evaluate the following expression when a = 2 and b = 5. (3ab)^b-4

OpenStudy (anonymous):

Just put the value you will get the solution.

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

Like (325)^5-4 ???

OpenStudy (anonymous):

yes like that.

OpenStudy (anonymous):

325 = 60

OpenStudy (anonymous):

30 its mistake.

OpenStudy (anonymous):

Wait, how did you get 325 = 60?

OpenStudy (anonymous):

its multiplication actually 3*2*5 = 30

OpenStudy (anonymous):

which is your question \[(3ab)^{b - 4} \\ or \\ (3ab)^b - 4\]

OpenStudy (anonymous):

So its (30)^5-4? Is that right? And then I just solve?

OpenStudy (anonymous):

Which of those is your question? I have asked above. First one you will get = 30 , second one 30^5 - 4 which is equivalent to \[3^5.10^5 - 4 = 243.10^5 - 4 = ??\]

OpenStudy (anonymous):

Sorry, I'm still really confused.

OpenStudy (anonymous):

No I am asking what is the exact question then I can explain you clearly.

OpenStudy (anonymous):

I just want to know what the answer is and how it is solved simply.

OpenStudy (anonymous):

Here 3ab is represent multiplication something like 2*3*5 = 30 or 2*2*2 = 8 Similarly here 3ab = 3*a*b and a = 2, b = 5 so 3*2*5 = 30. ^ it represent power so 30^5 = 30*30*30*30*30 as 30 = 3 * 10 so 30^5 = 3*10*3*10*3*10*3*10*3*10 = 3*3*3*3*3*10*10*10*10*10 = 243*100000 = 24300000

OpenStudy (anonymous):

So you answer is 243,00000?

OpenStudy (anonymous):

thats why I asked you whats the power is it (b - 4) or b.

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

if the question is something like this \[(3ab)^{b-4} = 30 \\ (3ab)^{b} - 4 = 24299996\]

OpenStudy (anonymous):